# Fundamental multiplication formulas

### v1

@@ 3.5. A geometric setting for $H^{\textsc{b}}_{{\boldsymbol{q}},1}$
Recall the Hecke algebra $H_{{\boldsymbol{q}},1}=H_{{\boldsymbol{q}},1}^{\textsc{b}}$ of type $B_{r}$ from §2. Let $H_{q,1}:=H_{{\boldsymbol{q}},1}|_{{\boldsymbol{q}}=q}:=H_{{\boldsymbol{q}},1}\otimes_{\mathcal{A}}\mathbb{Z}$ be the $\mathbb{Z}$-algebra obtained by specialising ${\boldsymbol{q}}$ to the prime power $q$. Then
@@ 3.6. A geometric setting for $S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)$
Let $S^{\textsc{b}}_{q,1}(n,r):=S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)|_{{\boldsymbol{q}}=q}=S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)\otimes_{\mathcal{A}}\mathbb{Z}$ be the $\mathbb{Z}$-algebra obtained by specialising ${\boldsymbol{q}}$ to the prime power $q$. Then
@@
Let $G(q)={\text{\rm O}}_{2r}(q)$, ${\mathcal{X}}=\mathcal{F}_{n,r}^{\jmath}$ and $\mathfrak{F}_{G(q)}({\mathcal{X}}\times{\mathcal{X}})=\mathfrak{F}_{G(q)}({\mathcal{X}}\times{\mathcal{X}},\mathbb{Z})$. Recall the standard basis $\{e_{A}|_{{\boldsymbol{q}}=q}\mid A\in\Xi_{N,2r}\}$ for $S^{\textsc{b}}_{q,1}(n,r)$ and the orbital basis $\{f_{\mathcal{O}}\}$ for $\mathfrak{F}_{G(q)}({\mathcal{X}}\times{\mathcal{X}})$
@@
4. A standard basis and geometric setting for $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$
We describe the standard (or double coset, or orbital) basis for the ${\boldsymbol{q}}$-Schur algebra of type $D$ according to its algebraic or geometric setting.
4.1. The algebraic definition of the standard basis
Recall the Hecke algebras $H_{{\boldsymbol{q}},1}$ of type $B_{r}$ and its subalgebra $\check{H}$ of type $D_{r}$. By restriction, every $H_{{\boldsymbol{q}},1}$-module $M$ is an $\check{H}$-module, denoted by $M|_{\check{H}}$. For any $\alpha\in\Lambda^{\textsc{d}}(n,r)$, let ${\check{x}}_{\alpha}=\sum_{w\in{\check{W}}_{\alpha}}T_{w}$.
@@ 4.1. The algebraic definition of the standard basis
$x_{\lambda}H_{{\boldsymbol{q}},1}|_{\check{H}}=\begin{cases}{\check{x}}_{\lambda^{+}}\check{H}\oplus{\check{x}}_{\lambda^{-}}\check{H},&\text{ if }\lambda_{n+1}=0;\\ {\check{x}}_{\lambda^{\bullet}}\check{H},&\text{ if }\lambda_{n+1}\neq 0.\end{cases}$
@@
Since $W=\check{W}\cup s_{r}\check{W}$ is a disjoint union, we have $H_{{\boldsymbol{q}},1}=\check{H}\oplus T_{r}\check{H}$. Thus,
$x_{\lambda}H_{{\boldsymbol{q}},1}=\begin{cases}x_{\lambda}\check{H}\oplus x_{\lambda}T_{r}\check{H},&\text{ if }\lambda_{n+1}=0;\\ x_{\lambda}\check{H},&\text{ if }\lambda_{n+1}\neq 0.\\ \end{cases}$
If $\lambda_{n+1}=0$, then $x_{\lambda}={\check{x}}_{\lambda^{+}}$ and $T_{r}x_{\lambda}T_{r}={\check{x}}_{\lambda^{-}}$ by Corollary 2.7(1)&(2). We have $x_{\lambda}\check{H}={\check{x}}_{\lambda^{+}}\check{H}$ and $x_{\lambda}T_{r}\check{H}=T_{r}{\check{x}}_{\lambda^{-}}\check{H}\cong{\check{x}}_{\lambda^{-}}\check{H}$. If $\lambda_{n+1}\neq 0$, by Corollary 2.7(3), we have either $x_{\lambda}={\check{x}}_{\lambda^{\bullet}}$ or $x_{\lambda}=(1+T_{r}){\check{x}}_{\lambda^{\bullet}}$. In both cases, we have $x_{\lambda}H_{{\boldsymbol{q}},1}|_{\check{H}}\cong{\check{x}}_{\lambda^{\bullet}}\check{H}$. ∎
@@
There are ${\mathcal{A}}$-algebra isomorphism $S_{\boldsymbol{q}}^{\kappa}(n,r)\cong S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. By base change, it extends to a ${\mathcal{Z}}$-algebra isomorphism ${\mathcal{S}}_{\upsilon}^{\kappa}(n,r)\cong{\mathcal{S}}_{\upsilon}^{\textsc{d}}(n,r).$
@@
$g:\bigoplus_{\lambda\in\Lambda(n+1,r)}x_{\lambda}H_{{\boldsymbol{q}},1}|_{\check{H}}\longrightarrow\bigoplus_{\alpha\in\Lambda^{\textsc{d}}(n,r)}{\check{x}}_{\alpha}\check{H}.$
@@
Similar to Corollary 2.5, the above isomorphism allows us to define the standard basis for $S_{\boldsymbol{q}}^{\kappa}(n,r)$.
@@
The set $\{\phi_{\alpha,\beta}^{d}\mid(\alpha,d,\beta)\in\check{\mathcal{D}}(n,r)\}$ forms a basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. We call this basis the standard basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. Moreover, the idempotents $1_{\alpha}$ considered in Lemma 2.11 take the form $1_{\alpha}=\phi_{\alpha,\alpha}^{1}$.
@@ 4.2. Parametrising ${\text{\rm SO}}_{2r}(q)$-orbits
We now construct the standard basis for $S_{\boldsymbol{q}}^{\kappa}(n,r)$ defined in (2.1.1) via the isomorphism in Proposition 3.10.
Let ${\check{G}}={\check{G}}(q)={\text{\rm SO}}_{2r}(q)$ (and $G=G(q)={\text{\rm O}}_{2r}(q)$). We start with parametrising ${\check{G}}$-orbits in the complete flag variety ${\mathcal{B}}_{r}$ introduced in Section 3. Let $\mathsf{F}=\mathsf{F}^{s}$ be the standard complete flag with $\mathsf{F}_{i}=\langle e_{1},\ldots,e_{i}\rangle$ and let
@@
To parametrising ${\check{G}}(q)$-orbits in the $n$-step isotropic (partial) flag variety ${\mathcal{X}}:=\mathcal{F}_{n,2r}^{\jmath}$, we use the $G(q)$-orbits. Let ${\mathcal{X}}/G(q)$ denote the set of all $G(q)$-orbits. Then (3.2.1) implies a bijection
@@
Any orbit ${\mathcal{O}}_{\lambda}=G(q).F$ splits into two ${\check{G}}(q)$-orbits if and only if $\lambda_{n+1}=0$.
@@
Hence, $G(q).F={\check{G}}(q).F\sqcup{\check{G}}(q)\dot{s}_{r}.F\iff\lambda_{n+1}=0$. The lemma follows ∎
@@
We now describe the decomposition of ${\mathcal{X}}\times{\mathcal{X}}$ into ${\check{G}}(q)$-orbits..
If $G(q).(F,F^{{}^{\prime}})$ is an $G(q)$-orbit in ${\mathcal{X}}\times{\mathcal{X}}$, then it can be written as $G(q).(F,F^{{}^{\prime}})={\check{G}}(q).(F,F^{{}^{\prime}})\cup{\check{G}}(q)g.(F,F^{{}^{\prime}})$ for some $g\in G(q)\setminus{\check{G}}(q)$. Since the map ${\mathfrak{m}}$ is a bijection (Corollary 5.7), we shall denote a $G(q)$-orbit by $\mathcal{O}_{A}:=G(q).(F,F^{{}^{\prime}})$ if ${\mathfrak{m}}(F,F^{\prime})=A$, for some $A\in\Xi_{N,2r}$.
@@
For $A\in\Xi_{N,2r}$, we have a $G(q)$-orbit $\mathcal{O}_{A}$ splits into two ${\check{G}}(q)$-orbits if and only if $a_{n+1,n+1}=0$.
@@
$\Phi:\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath}\longrightarrow\mathcal{F}_{n(N+1),r}^{\jmath},\ ({F},F^{{}^{\prime}})\longmapsto F_{\!\centerdot\centerdot}=(F_{i,j})_{1\leq i,j\leq N}.$ (4.5.1)
@@
    ${}^{+}\mathcal{O}_{A}:=\Phi_{A}^{-1}\left(\mathcal{F}_{n(N+1),d}^{\jmath}(A)^{+}\right)\text{ and }{}^{-}\mathcal{O}_{A}:=\Phi_{A}^{-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A)^{-}\right).$
@@
$\mathcal{O}_{A}^{+}:=\Phi_{A}^{\prime-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t})^{-}\right)\text{ and }\mathcal{O}_{A}^{-}:=\Phi_{A}^{\prime-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t})^{-}\right).$
@@
(1) For $A\in\overset{\circ}{\Xi}$ and $\epsilon_{1},\epsilon_{2}\in\{+,-\}$, we use ${}^{\epsilon_{1}}A^{\epsilon_{2}}$ to label the ${\check{G}}$-orbit ${}^{\epsilon_{1}}\mathcal{O}_{A}^{\epsilon_{2}}$ and write
@@
$\overset{\star\circ*}{\Xi}_{\scriptscriptstyle{-}}$ :=$:=\{{{}^{+}{A}^{-}},{{}^{-}{A}^{+}}\mid A\in\overset{\star\circ*}{\Xi}\text{ and }\text{\rm sgn}(A^{\text{\normalsize$\llcorner$}})=-\}.$
@@
Consider the diagonal embedding ${\mathcal{X}}\to{\mathcal{X}}\times{\mathcal{X}},F\mapsto(F,F)$. This map is ${\check{G}}(q)$-equivariant as well as $G(q)$-equivariant. Thus, the parametrisation in (4.4.1) is compatible with that in Definition 4.6. In other words, it induces injective maps
$\Lambda(n,r)\longrightarrow\Xi_{N,2r},\lambda\longmapsto D_{\lambda}:=\text{diag}(\lambda)\;\;\text{ and }\;\;\Lambda^{\textsc{d}}(n,r)\longrightarrow\check{\Xi}(n,r)$
@@
4.3. The geometric definition of the standard basis
Recall from Lemma 2.2 the bijective $\mathfrak{d}(\lambda,d,\mu)=(|R^{\lambda}_{i}\cap dR^{\mu}_{j}|)_{ij}$. If $\mathfrak{d}(\lambda,d,\mu)=A$, then we write $\mathfrak{d}^{-1}(A)=(\lambda,d_{A},\mu)$, where $\lambda=\text{\rm ro}(A)$, $\mu=\text{\rm co}(A)$. Note that $d_{A}\in{\check{W}}$ if and only if the entries of the upper right corner matrix $A^{\text{\normalsize$\llcorner$}}$ sum to a even number.

### v2

@@ 3.5. A geometric setting for $H^{\textsc{b}}_{{\boldsymbol{q}},1}$
Recall the Hecke algebra $H_{{\boldsymbol{q}},1}=H_{{\boldsymbol{q}},1}^{\textsc{b}}$ of type $B_{r}$ from §2. Let $H_{q,1}:=H_{{\boldsymbol{q}},1}|_{{\boldsymbol{q}}=q}:=H_{{\boldsymbol{q}},1}\otimes_{\mathcal{A}}\mathbb{Z}$ be the $\mathbb{Z}$-algebra obtained by specializing ${\boldsymbol{q}}$ to the prime power $q$. Then
@@ 3.6. A geometric setting for $S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)$
Let $S^{\textsc{b}}_{q,1}(n,r):=S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)|_{{\boldsymbol{q}}=q}=S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)\otimes_{\mathcal{A}}\mathbb{Z}$ be the $\mathbb{Z}$-algebra obtained by specializing ${\boldsymbol{q}}$ to the prime power $q$. Then
@@
Let $G(q)={\text{\rm O}}_{2r}(q)$, ${\mathcal{X}}=\mathcal{F}_{n,r}^{\jmath}$ and $\mathfrak{F}_{G(q)}({\mathcal{X}}\times{\mathcal{X}})=\mathfrak{F}_{G(q)}({\mathcal{X}}\times{\mathcal{X}},\mathbb{Z})$. Recall the natural basis $\{e_{A}|_{{\boldsymbol{q}}=q}\mid A\in\Xi_{N,2r}\}$ for $S^{\textsc{b}}_{q,1}(n,r)$ and the orbital basis $\{f_{\mathcal{O}}\}$ for $\mathfrak{F}_{G(q)}({\mathcal{X}}\times{\mathcal{X}})$
@@
4. A natural basis and geometric setting for $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$
We describe the natural (or double coset, or orbital) basis for the ${\boldsymbol{q}}$-Schur algebra of type $D$ according to its algebraic or geometric setting.
4.1. The algebraic definition of the natural basis
Recall the Hecke algebra $H=H_{{\boldsymbol{q}},1}$ of type $B_{r}$ and its subalgebra $\check{H}=\check{H}_{\boldsymbol{q}}$ of type $D_{r}$. By restriction, every $H$-module $M$ is an $\check{H}$-module, denoted by $M|_{\check{H}}$. For any $\alpha\in\Lambda^{\textsc{d}}(n,r)$, let ${\check{x}}_{\alpha}=\sum_{w\in{\check{W}}_{\alpha}}T_{w}$.
@@ 4.1. The algebraic definition of the natural basis
$x_{\lambda}H|_{\check{H}}\cong\begin{cases}{\check{x}}_{\lambda^{+}}\check{H}\oplus{\check{x}}_{\lambda^{-}}\check{H},&\text{ if }\lambda_{n+1}=0;\\ {\check{x}}_{\lambda^{\bullet}}\check{H},&\text{ if }\lambda_{n+1}\neq 0.\end{cases}$
@@
Since $W=\check{W}\cup s_{r}\check{W}$ is a disjoint union, we have $H=\check{H}\oplus T_{r}\check{H}$. Thus,
$x_{\lambda}H=\begin{cases}x_{\lambda}\check{H}\oplus x_{\lambda}T_{r}\check{H},&\text{ if }\lambda_{n+1}=0;\\ x_{\lambda}\check{H},&\text{ if }\lambda_{n+1}\neq 0.\\ \end{cases}$
If $\lambda_{n+1}=0$, then $x_{\lambda}={\check{x}}_{\lambda^{+}}$ and $T_{r}x_{\lambda}T_{r}={\check{x}}_{\lambda^{-}}$ by Corollary 2.7(1)&(2). We have $x_{\lambda}\check{H}={\check{x}}_{\lambda^{+}}\check{H}$ and $x_{\lambda}T_{r}\check{H}=T_{r}{\check{x}}_{\lambda^{-}}\check{H}\cong{\check{x}}_{\lambda^{-}}\check{H}$. If $\lambda_{n+1}\neq 0$, by Corollary 2.7(3), we have either $x_{\lambda}={\check{x}}_{\lambda^{\bullet}}$ or $x_{\lambda}=(1+T_{r}){\check{x}}_{\lambda^{\bullet}}$. In both cases, we have $x_{\lambda}H|_{\check{H}}\cong{\check{x}}_{\lambda^{\bullet}}\check{H}$. ∎
@@
There is an ${\mathcal{A}}$-algebra isomorphism $S_{\boldsymbol{q}}^{\kappa}(n,r)\cong S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. By base change, it extends to a ${\mathcal{Z}}$-algebra isomorphism ${\mathcal{S}}_{\upsilon}^{\kappa}(n,r)\cong{\mathcal{S}}_{\upsilon}^{\textsc{d}}(n,r).$
@@
$g:\bigoplus_{\lambda\in\Lambda(n+1,r)}x_{\lambda}H|_{\check{H}}\longrightarrow\bigoplus_{\alpha\in\Lambda^{\textsc{d}}(n,r)}{\check{x}}_{\alpha}\check{H}.$
@@
Similar to Corollary 2.5, the above isomorphism allows us to define the natural basis for $S_{\boldsymbol{q}}^{\kappa}(n,r)$.
@@
The set $\{\phi_{\alpha,\beta}^{d}\mid(\alpha,d,\beta)\in\check{\mathcal{D}}(n,r)\}$ forms a basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. We call this basis the natural basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. Moreover, the idempotents $1_{\alpha}$ considered in Lemma 2.11 take the form $1_{\alpha}=\phi_{\alpha,\alpha}^{1}$.
@@ 4.2. Parametrising ${\text{\rm SO}}_{2r}(q)$-orbits
We now construct the natural basis for $S_{\boldsymbol{q}}^{\kappa}(n,r)$ defined in (2.1.1) via the isomorphism in Proposition 3.10.
Let ${\check{G}}={\check{G}}(q)={\text{\rm SO}}_{2r}(q)$ (and $G=G(q)={\text{\rm O}}_{2r}(q)$). We start with parametrizing ${\check{G}}$-orbits in the complete flag variety ${\mathcal{B}}_{r}$ introduced in Section 3. Let $\mathsf{F}=\mathsf{F}^{s}$ be the standard complete flag with $\mathsf{F}_{i}=\langle e_{1},\ldots,e_{i}\rangle$ and let
@@
To parametrize ${\check{G}}(q)$-orbits in the $n$-step isotropic (partial) flag variety ${\mathcal{X}}:=\mathcal{F}_{n,2r}^{\jmath}$, we use the $G(q)$-orbits. Let ${\mathcal{X}}/G(q)$ denote the set of all $G(q)$-orbits. Then (3.2.1) implies a bijection
@@
An orbit ${\mathcal{O}}_{\lambda}=G(q).F$ splits into two ${\check{G}}(q)$-orbits if and only if $\lambda_{n+1}=0$.
@@
Hence, $G(q).F={\check{G}}(q).F\sqcup{\check{G}}(q)\dot{s}_{r}.F\iff\lambda_{n+1}=0$. The lemma follows. ∎
@@
We now describe the decomposition of ${\mathcal{X}}\times{\mathcal{X}}$ into ${\check{G}}(q)$-orbits.
If $G(q).(F,F^{{}^{\prime}})$ is a $G(q)$-orbit in ${\mathcal{X}}\times{\mathcal{X}}$, then it can be written as $G(q).(F,F^{{}^{\prime}})={\check{G}}(q).(F,F^{{}^{\prime}})\cup{\check{G}}(q)g.(F,F^{{}^{\prime}})$ for some $g\in G(q)\setminus{\check{G}}(q)$. Since the map ${\mathfrak{m}}$ is a bijection (Corollary 5.7), we shall denote a $G(q)$-orbit by $\mathcal{O}_{A}:=G(q).(F,F^{{}^{\prime}})$ if ${\mathfrak{m}}(F,F^{\prime})=A$, for some $A\in\Xi_{N,2r}$.
@@
Let $A=(a_{ij})\in\Xi_{N,2r}$. The $G(q)$-orbit $\mathcal{O}_{A}$ splits into two ${\check{G}}(q)$-orbits if and only if $a_{n+1,n+1}=0$.
@@
$\Phi:\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath}\longrightarrow\mathcal{F}_{n(N+1),r}^{\jmath},\ ({F},F^{{}^{\prime}})\longmapsto F_{\!\centerdot\centerdot}=(F,F^{\prime})_{\!\centerdot\centerdot}=(F_{i,j})_{1\leq i,j\leq N},$ (4.5.1)
@@
    ${}^{+}\mathcal{O}_{A}:=\Phi_{A}^{-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A)^{+}\right)\text{ and }{}^{-}\mathcal{O}_{A}:=\Phi_{A}^{-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A)^{-}\right).$
@@
$\mathcal{O}_{A}^{+}:=\Phi_{A}^{\prime-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t})^{+}\right)\text{ and }\mathcal{O}_{A}^{-}:=\Phi_{A}^{\prime-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t})^{-}\right).$
@@
(1) For $A\in\overset{\circ}{\Xi}$ and $\epsilon_{1},\epsilon_{2}\in\{+,-\}$, we use ${}^{\epsilon_{1}}A^{\epsilon_{2}}$ to label the ${\check{G}}(q)$-orbit ${}^{\epsilon_{1}}\mathcal{O}_{A}^{\epsilon_{2}}$ and write
@@
$\overset{\star\circ*}{\Xi}_{\scriptscriptstyle{-}}$ :=$:=\{{{}^{+}{A}^{-}},{{}^{-}{A}^{+}}\mid A\in\overset{\star\circ*}{\Xi}\text{ and }\text{\rm sgn}(A^{\text{\normalsize$\llcorner$}})=-\},$
@@
Consider the diagonal embedding ${\mathcal{X}}\to{\mathcal{X}}\times{\mathcal{X}},F\mapsto(F,F)$. This map is ${\check{G}}(q)$-equivariant as well as $G(q)$-equivariant. Thus, the parametrization in (4.4.1) is compatible with that in Definition 4.6. In other words, it induces injective maps
$\Lambda(n,r)\longrightarrow\Xi_{N,2r},\lambda\longmapsto D_{\lambda}:=\text{diag}(\lambda)\;\;\text{ and }\;\;\Lambda^{\textsc{d}}(n,r)\longrightarrow\check{\Xi}(n,r).$
@@
4.3. The geometric definition of the natural basis
Recall from Lemma 2.2 the bijective $\mathfrak{d}(\lambda,d,\mu)=(|R^{\lambda}_{i}\cap dR^{\mu}_{j}|)_{ij}$. If $\mathfrak{d}(\lambda,d,\mu)=A$, then we write $\mathfrak{d}^{-1}(A)=(\lambda,d_{A},\mu)$, where $\lambda=\text{\rm ro}(A)$, $\mu=\text{\rm co}(A)$. Note that $d_{A}\in{\check{W}}$ if and only if the entries of the upper right corner matrix $A^{\text{\normalsize$\llcorner$}}$ sum to an even number.